OpenStudy (anonymous):

A projectile is fired from the top of a building. It's initial position is (0m, 100m) and its initial velocities in the x and y directions are <32.9, 64.5>. Determine the time to reach the highest point, the coordinates of the highest point, time to reach the ground, and the coordinates of the point where the projectile hits the ground. I know it's long but I need help!

OpenStudy (amistre64):

are those velocities in meters per second?

OpenStudy (anonymous):

they are!

OpenStudy (amistre64):

\[h(t)=-\frac{g}{2}t^2+V_ot+H_o\]

OpenStudy (amistre64):

the highest point is determined by the "vertex" of this parabola, namely when:\[t=\frac{V_o}{g}\]

OpenStudy (amistre64):

those Vo are teh velocity in the up direction since height is an upward direction

OpenStudy (amistre64):

we find the time that the height is zero (it lands) by the quadratic formula:\[\frac{-V_o\pm\sqrt{{V_o}^2+4(\frac{g}{2})(H_o)}}{-g}\]

OpenStudy (anonymous):

oh that all makes sense! you rock.

OpenStudy (amistre64):

the distance outward then is a product of the x velocity and the time it takes to land.\[d(t)=V_{ox}\ t\] where t is the value from that quadratic formula form above

OpenStudy (anonymous):

thanks so much!